# Divergences in the hadronic light-by-light amplitude of the holographic soft-wall model

### v1

@@ VI.3 V-QCD
This is augmented by a scalar-extended Chern-Simons action generalizing further the one given above [66].
@@
times an oscillating function of period $\sqrt{E_{n}}$ with an average height that has a power-law behavior in $E_{n}$. The parameter $a_{\mathrm{IR}}$ parametrizes the tachyon potential $V_{f}$ in the DBI action, while $b$ appears in their scalar extended Chern-Simons term. The TFF will have exponential scaling behavior similar to (III.2) if $b-\frac{a_{\mathrm{IR}}}{2}<0$. Thus, in V-QCD, the issues discussed here can be avoided by taking $b>\frac{a_{\mathrm{IR}}}{2}$ (which is indeed the case for the parameter set in [65]).
@@ VII Conclusion and discussion
The hadronic light-by-light tensor is, however, divergent even when directly calculated from the 5-dimensional framework with bulk-to-bulk propagators. We therefore conclude that the conventional SW model as employed in [29, 30] is unsuitable for an evaluation of the hadronic light-by-light contribution to the muon g2g-2. Unlike the HW models studied previously, the SW model fails to provide a hadronic model where the MV-SDC can be satisfied. [^17] Ref. [30] reports that the MV-SDC was satisfied in the SW model. This will indeed be the case when the SW model is evaluated with a fixed cutoff on $z$. However, the numerical results for the HLbL amplitude as well as the $a_{\mu}$ contribution of the tower of resonances in the axial sector will depend on the cutoff, yielding infinity when the cutoff is removed. Even when used only as a model for the lowest excitations, the SW model does not appear to be in good shape. As noted already in [29], the SW model is in poor agreement for the ground-state pion contribution. Experimental and dispersive results for the $\pi^{0}$ TFF are not well reproduced, and $a_{\mu}^{\pi^{0}}$ is clearly overestimated when compared to dispersive and lattice results. On the other hand, the HW model introduced in [10] as well as variants thereof as studied in [57] agree remarkably well with phenomenological results. Similarly, the $a_{\mu}$ contribution from the $a_{1}$ axial-vector meson is larger in the SW model than in the HW model, whose prediction for the singly virtual TFF predicted by the latter for the $a_{1}$ has been found to agree surprisingly well with the recent dispersive results of [60].
@@ Acknowledgments
We would like to thank Luigi Cappiello and Harald Skarke for helpful discussions. This work has been supported by the Austrian Science Fund FWF, Grant-DOI 10.55776/PAT7221623.
@@ Appendix A Details on the WKB approximation
In this appendix we first recapitulate the WKB approximation and apply it to find the pseudoscalar and axial vector mode functions for high $m_{n}$ in the minimal soft-wall model described in the introduction.
@@ A.1 Pseudoscalar modes
In the following part, we will spell out some of the details omitted in the main part of the paper concerning the WKB approximation for the pseudoscalar sector. Axial vectors will be investigated in the next subsection. The Schrödinger equation for the pseudoscalar modes with $k^{2}(z)=E-V_{P}(z)$ is amenable to a WKB analysis for generic values of $z$. In this appendix, we will spell out in which regions the WKB approximation fails and which approximation to use instead. We will also find the points $z_{i}$ at which the solutions to the right and left are both valid and can therefore be matched at large $E_{n}$ Since the potential is bounded from below and the solution needs to be normalizable, we expect that solutions exist only for discrete values of $E=E_{n}$. The WKB approximation breaks down when $\frac{k^{\prime}}{k^{2}}\ll 1$ is violated. For any given fixed $z$, one may observe that $\frac{k^{\prime}}{k^{2}}\rightarrow 0$ as $E_{n}\rightarrow\infty$. For any large but fixed $E$, there are, however, always regions that are not described well by the WKB approximation. These are the regions around the two zeros $z_{l},z_{r}$ of $k^{2}(z)$ which at large $E_{n}$ are given approximately by as well as the very small-$z$ region where $\kappa^{2}\approx\frac{15}{4z^{2}}$ and therefore $\frac{\kappa^{\prime}}{\kappa^{2}}\sim\frac{1}{z^{4}}$. In these regions, the differential equation needs to be solved with different methods. If we pick $z_{1}(E_{n})$ in such a way that it approaches 0 in the limit, then we may always approximate $\kappa^{2}=E_{n}-V_{P}\sim E_{n}-\frac{15}{4z^{2}}$ in the whole range $z\in[0,z_{1}]$ at large $E_{n}$. With this form of $\kappa^{2}$, the analytic form of the solution that is normalizable is (28). Clearly, it is also a good approximation at $z=z_{1}$. We however do not want to let $z_{1}(E_{n})$ go to zero too quickly. The point $z_{l}$ at which the WKB solution diverges goes to zero like $E_{n}^{-\frac{1}{2}}$ and $z_{1}$ should always be larger than $z_{l}$ in order to have a chance of the WKB approximation being valid at $z_{1}$. Computing $\frac{\kappa^{\prime}}{\kappa^{2}}$ at $z=z_{1}$ with the ansatz $z_{1}=z_{l}\left(\frac{E_{n}}{c^{2}}\right)^{a_{1}}$ and demanding it to go to zero as $E_{n}\rightarrow\infty$ reveals
@@ A.2 Axial vector modes and nonzero quark mass
Since the potential for the axial vector mesons is qualitatively very similar to the pseudoscalar potential, the above analysis applies almost immediately to this case as well. The point $z_{l}$ is now at $\sqrt{\frac{3}{4E_{n}}}$ and in region I the solution changes to
@@ Appendix B Transition form factors of HW models in WKB approximation
Appendix C WKB approximation to the axial bulk-to-bulk propagator in the SW model
@@
Appendix D Nonuniform convergence of mode sums
As we will now argue, the apparent contradiction is caused by performing an illegal exchange of limits.
@@ Appendix D Nonuniform convergence of mode sums
$\langle A^{\perp}_{n},\mathcal{A}^{\perp}\rangle=-\frac{F_{n}^{A}}{Q^{2}+m_{n}^{2}}.$ (155)
@@
$\mathcal{A}^{\perp}_{\text{sum}}=-\sum_{n}\frac{F^{A}_{n}}{Q^{2}+m_{A}^{2}}A^{\perp}_{n}(z)$ (156)
@@
In appendix A, the axial decay constant scaling is determined to be $F_{n}^{A}\sim C_{1}^{A}\sqrt{E_{n}}\sim E_{n}^{\frac{1}{8}}\sqrt{E_{n}}$. We may also approximate $Q^{2}+m_{n}^{2}\sim m_{n}^{2}=E_{n}$ at sufficiently high $E_{n}=m_{n}^{2}\sim n^{4/3}$.
@@
The convergence of $\xi_{\text{sum}}$ is also uniform on any compact interval excluding zero (a small region $0,\delta)$ around $z=0$ is excluded as explained in footnote 19). We recapitulate some aspects of uniform convergence in Appendix E.
@@
$h_{N}(z)=\xi(z)-\sum_{n=0}^{N-1}\frac{-F_{n}^{A}}{Q^{2}+m_{n}^{2}}\xi_{n}(z)=\sum_{n=N}^{\infty}\frac{-F_{n}^{A}}{Q^{2}+m_{n}^{2}}\xi_{n}(z)$ (166)
@@
It may seem puzzling that $\xi_{\text{sum}}$ converges uniformly on $(\delta,\infty)$, but $\mathcal{A}^{\perp}_{\text{sum}}$ does not since the individual summands just differ by multiplication by $\frac{1}{\sqrt{\rho}}=e^{+\frac{c^{2}z^{2}}{2}}\sqrt{z}$. We explain how this can come about with a simple example in appendix E.
@@
Appendix E Uniform convergence
@@
In Appendix D, we claimed that $\xi_{\text{sum}}$ converges uniformly on $\delta,\infty]$, whilst $\mathcal{A}^{\perp}_{\text{sum}}$ does not. Here we would like to illustrate that this is possible by a simple example. Pick a sequence of functions $f_{n}(z)=\frac{1}{n}\theta\left(z-(z_{n}-\Delta)\right)\theta\left(-z+(z_{n}+\Delta)\right)$with $z_{n}$ a monotonic sequence that goes to infinity. This sequence describes rectangular bumps of fixed length that move to the right and whose height decreases as $n$ increases. Clearly $f_{n}$ converges to 0 both pointwise and uniformly since $h_{n}$ is a bounded function whose bound goes to zero as $n\rightarrow\infty$.

### v2

@@ VI.3 V-QCD
The potential $V_{f}$ depends on the tachyon in the following way
$V_{f}=V_{f}^{0}(\lambda)e^{-a_{IR}\,\tau_{0}^{2}}.$ (116)
The quadratic action is augmented by a scalar-extended Chern-Simons action generalizing further the one given above [66]. A parameter $b$ is added which introduces exponential damping for the Chern-Simons term of the form
$S_{CS}\sim\int e^{-b\,\tau_{0}^{2}}\times\left(\text{5-forms}\right).$ (117)
@@
times an oscillating function of period $\sqrt{E_{n}}$ with an average height that has a power-law behavior in $E_{n}$. The parameter $a_{\mathrm{IR}}$ parametrizes the tachyon potential $V_{f}$ in the DBI action, while $b$ appears in their scalar-extended Chern-Simons term. The TFF will have exponential scaling behavior similar to (III.2) if $b-\frac{a_{\mathrm{IR}}}{2}<0$.
In V-QCD one typically already chooses $b>a$ to obtain regular background solutions in the case of a nonzero $\theta$-angle [67, 68]. The issues of convergence discussed in this paper are thus avoided.
@@ VII Conclusion and discussion
The hadronic light-by-light tensor is, however, divergent even when directly calculated from the 5-dimensional framework with bulk-to-bulk propagators. We therefore conclude that the conventional SW model as employed in [29, 30] is unsuitable for an evaluation of the hadronic light-by-light contribution to the muon g2g-2. Unlike the HW models studied previously, the SW model fails to provide a hadronic model where the MV-SDC can be satisfied. [^17] Ref. [30] reports that the MV-SDC was satisfied in the SW model. This will indeed be the case when the SW model is evaluated with a fixed cutoff on $z$. However, the numerical results for the HLbL amplitude as well as the $a_{\mu}$ contribution of the tower of resonances in the axial sector will depend on the cutoff, yielding infinity when the cutoff is removed. Even when used only as a model for the lowest excitations, the SW model does not appear to be in good shape. As noted already in [29], the SW model is in poor agreement for the ground-state pion contribution. Experimental and dispersive results for the $\pi^{0}$ TFF are not well reproduced, and $a_{\mu}^{\pi^{0}}$ is clearly overestimated when compared to dispersive and lattice results. On the other hand, the HW model introduced in [10] as well as variants thereof as studied in [57] agree remarkably well with phenomenological results. Similarly, the $a_{\mu}$ contribution from the $a_{1}$ axial-vector meson is larger in the SW model than in the HW model, whose prediction for the singly virtual TFF predicted by the latter for the $a_{1}$ has been found to agree surprisingly well with the recent dispersive results of [18, 60, 23].
@@ Acknowledgments
We would like to thank Luigi Cappiello, Matti Järvinen, Elias Kiritsis, and Harald Skarke for helpful discussions. This work has been supported by the Austrian Science Fund FWF, Grant-DOI 10.55776/PAT7221623.
@@ Appendix A Details on the WKB approximation
In this appendix we recapitulate the WKB approximation and use it to find the pseudoscalar and axial vector mode functions for high masses in the minimal soft-wall model described in the introduction.
@@ A.1 Pseudoscalar modes
In the following part, we spell out some of the details omitted in the main part of the paper concerning the WKB approximation for the pseudoscalar sector. Axial vectors will be investigated in the next subsection. The Schrödinger equation for the pseudoscalar modes with $k^{2}(z)=E-V_{P}(z)$ is amenable to a WKB analysis for generic values of $z$. In this appendix, we will spell out in which regions the WKB approximation fails and which approximation to use instead. We will also find the points $z_{i}$ at which the solutions to the right and left are both valid and can therefore be matched at large $E_{n}$.
Since the potential is bounded from below and the solution needs to be normalizable, we expect that solutions exist only for discrete values of $E=E_{n}$. The WKB approximation breaks down when $\frac{k^{\prime}}{k^{2}}\ll 1$ is violated. For any given fixed $z$, one may observe that $\frac{k^{\prime}}{k^{2}}\rightarrow 0$ as $E_{n}\rightarrow\infty$. For any large but fixed $E$, there are, however, always regions that are not described well by the WKB approximation. These are the regions around the two zeros $z_{l},z_{r}$ of $k^{2}(z)$ which at large $E_{n}$ are given approximately by
$z_{l}$ $=\sqrt{\frac{15}{4E_{n}}},$
$z_{r}$ $=\bigg(\frac{E_{n}}{g_{5}^{2}\sigma^{2}}\bigg)^{\frac{1}{4}},$
as well as the very small-$z$ region where $\kappa^{2}\approx\frac{15}{4z^{2}}$ and therefore $\frac{\kappa^{\prime}}{\kappa^{2}}\sim\frac{1}{z^{4}}$. In these regions, the differential equation needs to be solved with different methods. If we pick $z_{1}(E_{n})$ in such a way that it approaches 0 in the limit, then we may always approximate $\kappa^{2}=E_{n}-V_{P}\sim E_{n}-\frac{15}{4z^{2}}$ in the whole range $z\in[0,z_{1}]$ at large $E_{n}$. With this form of $\kappa^{2}$, the analytic form of the solution that is normalizable is (28). Clearly, it is also a good approximation at $z=z_{1}$. We however do not want to let $z_{1}(E_{n})$ go to zero too quickly. The point $z_{l}$ at which the WKB solution diverges goes to zero like $E_{n}^{-\frac{1}{2}}$ and $z_{1}$ should always be larger than $z_{l}$ in order to have a chance of the WKB approximation being valid at $z_{1}$. Computing $\frac{\kappa^{\prime}}{\kappa^{2}}$ at $z=z_{1}$ with the ansatz $z_{1}=z_{l}\left(\frac{E_{n}}{c^{2}}\right)^{a_{1}}$ and demanding it to go to zero as $E_{n}\rightarrow\infty$ reveals
@@ A.2 Axial vector modes and nonzero quark mass
Since the potential for the axial vector mesons is qualitatively very similar to the pseudoscalar potential, the above analysis applies almost immediately to this case as well. The point $z_{l}$ is now approximately at
$z_{L}=\sqrt{\frac{3}{4E_{n}}}$ (139)
and in region I the solution changes to
@@ Appendix B Transition form factors of HW models in WKB approximation
Appendix C HLbL amplitude in the Witten-Sakai-Sugimoto model
In this appendix we investigate HLbL scattering and potential convergence issues in the Witten-Sakai-Sugimoto (WSS) model, which in contrast to the bottom-up models is based on a string-theoretic construction. While this model cannot be matched to the short-distance constraints of QCD and correspondingly is less interesting phenomenologically, regarding finiteness, the WSS model behaves essentially like a hard-wall model and no divergences are found.
As the tower of excited pions is absent in the WSS model, the contribution to the HLbL tensor in the pseudoscalar sector is given solely by the ground-state Goldstone bosons and is thus trivially finite. The model does however contain an infinite number of excited axial vector mesons, which we now analyze.
Both vector and axial vector meson modes are described by a $U(N_{f})$ gauge field $A$ on the flavor branes whose holographic direction [^19] We do not consider fluctuations on the $S^{4}$ part of the branes. is described by the coordinate $Z\in(-\infty,\infty)$. They are encoded in the even and odd parts of $A_{\mu}$, respectively. We therefore work on the half interval $Z\in[0,\infty]$. The vector and axial vector equations of motion are identical
$\partial_{Z}(K\partial_{Z}\mathcal{J})-K^{-1/3}\left(\frac{Q}{M_{\mathrm{KK}}}\right)^{2}\mathcal{J}=0,$ (154)
$\partial_{Z}(K\partial_{Z}\psi_{A})+K^{-1/3}\left(\frac{M^{A}}{M_{\mathrm{KK}}}\right)^{2}\psi_{A}=0,$ (155)
where $K=1+Z^{2}$, but differ in the IR boundary conditions
$\mathcal{J}^{\prime}(0,Q)=0,$
$\psi_{A}(0)=0.$ (156)
To apply the WKB machinery we transform the equations of motion into the form of a Schrödinger equation using the variable transform
$\frac{\partial x}{\partial Z}=K^{-2/3},$ (157)
which is solved by
$x(Z)=Z\,{}_{2}F_{1}(\frac{1}{2},\frac{2}{3};\frac{3}{2},-Z^{2}).$ (158)
This transformation maps the non-compact $Z$ interval to a compact interval $x\in[0,x_{f}]$ with $x_{f}=x(Z=\infty)=\frac{1}{2}B(\frac{1}{2},\frac{1}{6})$ and $B(a,b)$ the Euler beta function. With respect to these coordinates the IR boundary is at $x=0$ and the UV boundary at $x=x_{f}$. We then define $\xi$ via
$\psi_{A}=\frac{1}{\sqrt{\rho}}\xi,\qquad\rho=K^{1/3},$ (159)
and similarly for the vector sector, and in addition transform variables one more time to
$z=z_{0}-x,$ (160)
where $z_{0}=x_{f}$ to arrive at
$-\partial_{z}^{2}\xi+V_{\mathrm{WSS}}\,\xi=E\,\xi,$ (161)
with $E=(M^{A}/M_{\mathrm{KK}})^{2}$ or $E=-(Q/M_{\mathrm{KK}})^{2}$. Just like in HW models the asymptotic boundary is now at $z=0$ and the IR hard wall is at $z=z_{0}$. The Schrödinger potential $V_{\mathrm{WSS}}$ is given by
$V_{\mathrm{WSS}}=\frac{1}{3}\frac{1+\frac{2}{3}Z^{2}}{(1+Z^{2})^{2/3}},$ (162)
where $Z=Z(z)$ is the inverse of the composition of the two coordinate transformations (158) and (160). At small $z$, we have
$V_{\mathrm{WSS}}\sim\frac{2}{z^{2}},$ (163)
while close to $z=z_{0}$,
$V_{\mathrm{WSS}}\sim\frac{1}{3}.$ (164)
Similar to AdS models, the potential is singular near $z=0$, albeit with a different coefficient [^20] This coefficient is determined by the behavior of $\rho$. The WSS model obtains a different coefficient since it is secretly dual to a 5-dimensional field theory., and regular everywhere else. One may thus solve for mode functions, bulk-to-boundary and bulk-to-bulk propagators using the WKB approximation as before. Instead of proceeding in this direction, we now prove that the HLbL scattering amplitude is finite in the WSS model.
The axial-vector contributions to the coefficient functions $\hat{\Pi}_{i}$ of the HLBL tensor are all of the form
$\iint_{0}^{z_{0}}dzdz^{\prime}\widetilde{\mathcal{J}}(z,Q_{1})\mathcal{J}(z,Q_{2})\widetilde{\mathcal{J}}(z^{\prime},Q_{3})G_{A}(z,z^{\prime},P^{2}),$ (165)
where $G_{A}$ is the axial bulk-to-bulk propagator and $\widetilde{\mathcal{J}}(z,Q)={\partial_{z}\mathcal{J}(z,Q)}/{q^{2}}$, for some momenta $Q_{i},P$. The integrand is a continuous function of $z,z^{\prime}$ everywhere except near $z=0$ and $z^{\prime}=0$ because $\mathcal{J},\widetilde{\mathcal{J}},G_{A}$ are solutions to linear differential equations [^21] The axial bulk-to-bulk propagator has a discontinuous first derivative at $z=z^{\prime}$, but is itself continuous. whose coefficients are smooth except near $z=0$. Near $z=0$, we can determine the behavior of the solutions by using (163). The most general solution of (154) at small $z$ is given by
$c_{1}(Q)\left(\cosh(Qz)-Qz\,\sinh(Qz)\right)$
$+$ $c_{2}(Q)\left({\sinh(Qz)}-Qz\,\cosh(Qz)\right)$
$\equiv\mathcal{I}_{1}(z)+\mathcal{I}_{2}(z),$ (166)
which is regular at $z=0$. The vector bulk-to-boundary has $c_{1}(Q)=1$ and $c_{2}(Q)$ is fixed by matching to the correct IR solution. The axial bulk-to-bulk propagator $G_{A}$ is proportional to
$G_{A}\propto\mathcal{I}_{1}(z)\mathcal{I}_{2}(z^{\prime})\theta(z^{\prime}-z)+\mathcal{I}_{2}(z)\mathcal{I}_{1}(z^{\prime})\theta(z-z^{\prime}),$ (167)
and thus $\mathcal{J},\widetilde{\mathcal{J}},G_{A}$ are all continuous near $z=0$ and $z^{\prime}=0$. Thus the HLbL expressions (165) are integrals of continuous functions over a compact region and therefore finite. The divergences found in the soft-wall models are not present in the WSS model.
Appendix D WKB approximation to the axial bulk-to-bulk propagator in the SW model
@@
Appendix E Nonuniform convergence of mode sums
In section IV we found two different results for the VVA correlator, specifically the coefficient function $\bar{\mathcal{W}}_{2}(q^{2})$, depending on how the computation was performed. When evaluating the 5d Witten diagram using vector and axial vector bulk to boundary propagators a well defined and finite $\bar{\mathcal{W}}_{2}(q^{2})$ was found, while when reconstructing it dispersively using axial TFFs a divergence was encountered in the infinite sum over resonances. As we will now argue, this apparent contradiction arises from an illegal exchange of limits.
@@ Appendix E Nonuniform convergence of mode sums
$\langle A^{\perp}_{n},\mathcal{A}^{\perp}\rangle=-\frac{F_{n}^{A}}{Q^{2}+(M_{n}^{A})^{2}}.$ (172)
@@
$\mathcal{A}^{\perp}_{\text{sum}}=-\sum_{n}\frac{F^{A}_{n}}{Q^{2}+(M_{n}^{A})^{2}}A^{\perp}_{n}(z)$ (173)
@@
In appendix A, the axial decay constant scaling is determined to be $F_{n}^{A}\sim C_{1}^{A}\sqrt{E_{n}}\sim E_{n}^{\frac{1}{8}}\sqrt{E_{n}}$. We may also approximate $Q^{2}+(M_{n}^{A})^{2}\sim(M_{n}^{A})^{2}=E_{n}$ at sufficiently high $E_{n}=m_{n}^{2}\sim n^{4/3}$.
@@
The convergence of $\xi_{\text{sum}}$ is also uniform on any compact interval excluding zero (a small region $0,\delta)$ around $z=0$ is excluded as explained in footnote 22). We recapitulate some aspects of uniform convergence in Appendix F.
@@
$h_{N}(z)$ $=\xi(z)-\sum_{n=0}^{N-1}\frac{-F_{n}^{A}}{Q^{2}+(M_{n}^{A})^{2}}\xi_{n}(z)$
$=\sum_{n=N}^{\infty}\frac{-F_{n}^{A}}{Q^{2}+(M_{n}^{A})^{2}}\xi_{n}(z)$ (183)
@@
It may seem puzzling that $\xi_{\text{sum}}$ converges uniformly on $(\delta,\infty)$, but $\mathcal{A}^{\perp}_{\text{sum}}$ does not since the individual summands just differ by multiplication by $\frac{1}{\sqrt{\rho}}=e^{+\frac{c^{2}z^{2}}{2}}\sqrt{z}$. We explain how this can come about with a simple example in appendix F.
@@
Appendix F Uniform convergence
@@
In Appendix E, we claimed that $\xi_{\text{sum}}$ converges uniformly on $\delta,\infty]$, whilst $\mathcal{A}^{\perp}_{\text{sum}}$ does not. Here we would like to illustrate that this is possible by a simple example. Pick a sequence of functions $f_{n}(z)=\frac{1}{n}\theta\left(z-(z_{n}-\Delta)\right)\theta\left(-z+(z_{n}+\Delta)\right)$with $z_{n}$ a monotonic sequence that goes to infinity. This sequence describes rectangular bumps of fixed length that move to the right and whose height decreases as $n$ increases. Clearly $f_{n}$ converges to 0 both pointwise and uniformly since $h_{n}$ is a bounded function whose bound goes to zero as $n\rightarrow\infty$.